# Neural Networks for Fast Optimisation in Model Predictive Control: A Review

### v2

@@ 7.1 Problem types and the lack of benchmarks
Perhaps the more meaningful information pertaining to training time, is how it affects the performance of the resulting approximation. Figure 5d shows this relationship as reported by Li et al. Li et al. (2022) in their case studies. Data in the plot was generated for a neural approximation and control problem of fixed size, meaning that the plant and prediction horizon of the original controller remained constant for all experiments. Only the amount of data used to train the network was varied. As expected, increasing the amount of training data increases the required training time, but as Li et al. Li et al. (2022) noted, it also tends to increase performance. In the plot performance is expressed as “average cost value”. This metric is the mean of the objective function in the original MPC optimisation problem evaluated for every test point with the values produced by the NN approximation. Thus, better performance corresponds to a lower average cost. The linear regression of the average cost data in the plot confirms that performance does tend to increase with increasing amounts of training data. The gradient of the fit is small and its quality is low $(R^{2}=0.405)$ indicating that this effect is weak. Nevertheless, it is possible that not all approximation frameworks exhibit this relationship, or that they present different relationships with possibly stronger correlations. For practical applications knowing this information before applying a framework would be very useful. Another important point is that the framework by Li et al. Li et al. (2022) does not involve generating training data by repeatedly solving MPC problems. Instead, the method is an optimise and train approach where the MPC problems are solved at the same time as the NN approximation is trained, all as part of a unified problem. The time involved in data generation for other approaches is non-negligible. For instance, Chen et al. Chen et al. (2022) report up to $5.2\;hr$ for their largest case study despite using an efficient geometric walk approach. Therefore, in addition to measuring all aforementioned scaling properties, an adequate benchmarking tool should also aim to measure total approximation time, including both data generation and training time.
@@
Having discussed some of the comparisons that can be made when more rigorous testing is involved, we now make an effort to define what a truly informative benchmark should look like. First, the proposed benchmark should consider a wide range of problem types and dimensions for all versions of MPC, chosen with the intent of assessing scalability and applicability beyond specific cases. With such a suite of problems, the scaling relationships in Figures 5a and 5b could be defined for all frameworks. However, unlike some of the data presented in the figures, experiments should be conducted by varying only one variable at a time and conducting multiple experiments. For instance, the benchmark could measure how the required network size scales with plant size for horizons of fixed short, medium, and long lengths. Similarly, the scaling with horizon length could be measured for small, medium, and large plants. Such experiments should also involve a diverse combination of problem types and SISO and MIMO plants, so that the identified trends can be assumed to hold in general. In terms of computation time, each experiment should measure the time required to generate training data, training time, and the factor by which the resulting approximation speeds up the original controller. To facilitate fair comparisons between frameworks, these metrics should be recorded alongside detailed descriptions of the hardware and software used for the tests, including details of the implementation of the original MPC controller. Regarding performance, all tests should measure metrics that quantify the compromises involved by using the resulting approximation instead of the original MPC controller. Li et al. Li et al. (2022) provided good examples of such metrics including the “average cost value” reported in Figure 5d, as well as, the maximum and average constraint violation encountered during tests. However, since average cost value is problem specific, it would make more sense to normalise this value by dividing the average cost by the average cost of the original MPC controller. For a perfect approximation of the controller, this performance index would be 1. For imperfect approximations, it would be greater than 1 and become larger as the quality of the approximation decreases.
@@ 7.2 Areas for future research
We start our discussion of the future of the field by summarising the gaps that have become immediately apparent in reviewing all the works we have covered. This summary is a list of opportunities to make contributions in short to medium time frames. Regarding the development of new NNBO frameworks, there are no existing methods that address the general case of integer control variables. Existing studies have only treated the case of binary variables Löhr et al. (2019); Karg and Lucia (2018); Löhr et al. (2020) and in all cases only a single integer variable was considered. For NMPC only a few frameworks Parisini et al. (1998); Pin et al. (2010, 2013); Ahmed and Al-Dajani (1998) offer strict guarantees without relying on further optimisation or projection operations. However, as summarised in Table 2 these frameworks can only preserve select guarantees, they have only been demonstrated on small problems, and their achieved speedup factors has not been reported. For RMPC all frameworks with strict guarantees also rely on further optimisation or projection Daosud et al. (2019); Paulson and Mesbah (2020). Moreover, all research concerning strict guarantees and no guarantees has focused on msRMPC, while min-max RMPC and Tube RMPC remain unaddressed. Another topic that remains unexplored is the use of more advanced neural architectures to generate approximations. The studies we covered mostly use shallow networks or multi-layer perceptrons with standard neurons and conventional activation functions. Only a few studies explored the use of more involved ML tools, such as LSTM units Kumar et al. (2018) or autoencoders Nurbayeva et al. (2022). Many promising ML tools remain unused in the field, particularly those that can exploit temporal or sequential data and prior knowledge of the system. For instance, the use of transformers, which are best known for their use in natural language processing, has already been demonstrated in MPC for the purpose of modelling and their use in replacing the optimiser has also been suggested Park et al. (2023). Another promising option is the use of Physics Informed NNs (PINNs), which are designed to respect the underlying physics of the process they model. This type of network has been recently adapted for use in MPC as a plant replacement Antonelo et al. (2021). However, the use of PINNs as part of an NNBO framework has not been explored. Lastly, other application oriented contributions could also be made along the lines of developing open source libraries that implement approximation frameworks for general purpose use or other NNBO tools. As already discussed, the field could greatly benefit from the development of good benchmarking tools.
@@
Beyond the evident gaps, we believe that future breakthroughs in the field lie in the development of ties between control and learning theory. Karg and Lucia Karg and Lucia (2020b) and Fabiani and Gourlart Fabiani and Goulart (2022) pioneered this approach for LMPC. Their combined work resulted in a framework with strict optimality, stability, and recursive feasibility guarantees, which does not rely on post processing operations like projection or further optimisation. Instead, the result is a NN that can be used as a drop-in replacement for the original controller, just like for most pure regression approaches. Extending these results to NMPC and RMPC will require more theoretical work to create stronger links between control and learning theory. The lack of such ties has been noted more often in learning-based control research where it is especially difficult to guarantee the safety and robustness of control algorithms. Shi Shi (2023) treated this problem extensively noting that both the learning and control fields can greatly benefit from the development of unified frameworks that ease translation and encourage new discoveries. It is possible then that by studying efforts in translating control principles to learning principles, one might find ways to formulate new MPC approximation methods with strict guarantees. For instance, Lin et al. Lin et al. (2021) worked on proving the dynamic regret properties of LVMPC, thereby creating new translation tools between learning and control. None of the works reviewed here use or exploit the concept of regret. Alternatively, Shi et al. Shi et al. (2019) showed how spectral normalisation can be used to constrain the Lipschitz constant of a DNN and train nonlinear DNN-based controllers with strict stability guarantees. The resulting controller was purely learning-based, but it might be possible to adapt this novel training method for the purpose of approximating MPC with strict guarantees.

### v3

@@ 7.1 Problem types and the lack of benchmarks
Perhaps the more meaningful information pertaining to training time, is how it affects the performance of the resulting approximation. Figure 5d shows this relationship as reported by Li et al. (2022) in their case studies. Data in the plot was generated for a neural approximation and control problem of fixed size, meaning that the plant and prediction horizon of the original controller remained constant for all experiments. Only the amount of data used to train the network was varied. As expected, increasing the amount of training data increases the required training time, but as Li et al. (2022) noted, it also tends to increase performance. In the plot performance is expressed as “average cost value”. This metric is the mean of the objective function in the original MPC optimisation problem evaluated for every test point with the values produced by the NN approximation. Thus, better performance corresponds to a lower average cost. The linear regression of the average cost data in the plot confirms that performance does tend to increase with increasing amounts of training data. The gradient of the fit is small and its quality is low $(R^{2}=0.405)$ indicating that this effect is weak. Nevertheless, it is possible that not all approximation frameworks exhibit this relationship, or that they present different relationships with possibly stronger correlations. For practical applications knowing this information before applying a framework would be very useful. Another important point is that the framework by Li et al. (2022) does not involve generating training data by repeatedly solving MPC problems. Instead, the method is an optimise and train approach where the MPC problems are solved at the same time as the NN approximation is trained, all as part of a unified problem. The time involved in data generation for other approaches is non-negligible. For instance, Chen et al. (2022) reported up to $5.2\;hr$ for their largest case study despite using an efficient geometric walk approach. Therefore, in addition to measuring all aforementioned scaling properties, an adequate benchmarking tool should also aim to measure total approximation time, including both data generation and training time.
@@
Having discussed some of the comparisons that can be made when more rigorous testing is involved, we now make an effort to define what a truly informative benchmark should look like. First, the proposed benchmark should consider a wide range of problem types and dimensions for all versions of MPC, chosen with the intent of assessing scalability and applicability beyond specific cases. With such a suite of problems, the scaling relationships in Figures 5a and 5b could be defined for all frameworks. However, unlike some of the data presented in the figures, experiments should be conducted by varying only one variable at a time and conducting multiple experiments. For instance, the benchmark could measure how the required network size scales with plant size for horizons of fixed short, medium, and long lengths. Similarly, the scaling with horizon length could be measured for small, medium, and large plants. Such experiments should also involve a diverse combination of problem types and SISO and MIMO plants, so that the identified trends can be assumed to hold in general. In terms of computation time, each experiment should measure the time required to generate training data, training time, and the factor by which the resulting approximation speeds up the original controller. To facilitate fair comparisons between frameworks, these metrics should be recorded alongside detailed descriptions of the hardware and software used for the tests, including details of the implementation of the original MPC controller. Regarding performance, all tests should measure metrics that quantify the compromises involved by using the resulting approximation instead of the original MPC controller. Li et al. (2022) provided good examples of such metrics including the “average cost value” reported in Figure 5d, as well as, the maximum and average constraint violation encountered during tests. However, since average cost value is problem specific, it would make more sense to normalise this value by dividing the average cost by the average cost of the original MPC controller. For a perfect approximation of the controller, this performance index would be 1. For imperfect approximations, it would be greater than 1 and become larger as the quality of the approximation decreases.
@@ 7.2 Areas for future research
We start our discussion of the future of the field by summarising the gaps that have become immediately apparent in reviewing all the works we have covered. This summary is a list of opportunities to make contributions in short to medium time frames. Regarding the development of new NNBO frameworks, there are no existing methods that address the general case of integer control variables. Existing studies have only treated the case of binary variables (Löhr et al., 2019; Karg and Lucia, 2018; Löhr et al., 2020) and in all cases only a single integer variable was considered. For NMPC only a few frameworks in (Parisini et al., 1998; Pin et al., 2010, 2013; Ahmed and Al-Dajani, 1998) offer strict guarantees without relying on further optimisation or projection operations. However, as summarised in Table 2 these frameworks can only preserve select guarantees, they have only been demonstrated on small problems, and their achieved speedup factors has not been reported. For RMPC all frameworks with strict guarantees also rely on further optimisation or projection (Daosud et al., 2019; Paulson and Mesbah, 2020). Moreover, all research concerning strict guarantees and no guarantees has focused on msRMPC, while min-max RMPC and Tube RMPC remain unaddressed. Another topic that remains unexplored is the use of more advanced neural architectures to generate approximations. The studies we covered mostly use shallow networks or multi-layer perceptrons with standard neurons and conventional activation functions. Only a few studies explored the use of more involved ML tools, such as LSTM units (Kumar et al., 2018) or autoencoders (Nurbayeva et al., 2022). Many promising ML tools remain unused in the field, particularly those that can exploit temporal or sequential data and prior knowledge of the system. For instance, the use of transformers, which are best known for their use in natural language processing, has already been demonstrated in MPC for the purpose of modelling and their use in replacing the optimiser has also been suggested (Park et al., 2023). Another promising option is the use of Physics Informed NNs (PINNs), which are designed to respect the underlying physics of the process they model. This type of network has been recently adapted for use in MPC as a plant replacement (Antonelo et al., 2021). However, the use of PINNs as part of an NNBO framework has not been explored. Lastly, other application oriented contributions could also be made along the lines of developing open source libraries that implement approximation frameworks for general purpose use or other NNBO tools. As already discussed, the field could greatly benefit from the development of good benchmarking tools.
@@
Beyond the evident gaps, we believe that future breakthroughs in the field lie in the development of ties between control and learning theory. Karg and Lucia (2020b) and Fabiani and Goulart (2022) pioneered this approach for LMPC. Their combined work resulted in a framework with strict optimality, stability, and recursive feasibility guarantees, which does not rely on post processing operations like projection or further optimisation. Instead, the result is a NN that can be used as a drop-in replacement for the original controller, just like for most pure regression approaches. Extending these results to NMPC and RMPC will require more theoretical work to create stronger links between control and learning theory. The lack of such ties has been noted more often in learning-based control research where it is especially difficult to guarantee the safety and robustness of control algorithms. Shi (2023) treated this problem extensively noting that both the learning and control fields can greatly benefit from the development of unified frameworks that ease translation and encourage new discoveries. It is possible then that by studying efforts in translating control principles to learning principles, one might find ways to formulate new MPC approximation methods with strict guarantees. For instance, Lin et al. (2021) worked on proving the dynamic regret properties of LVMPC, thereby creating new translation tools between learning and control. None of the works reviewed here use or exploit the concept of regret. Alternatively, Shi et al. (2019) showed how spectral normalisation can be used to constrain the Lipschitz constant of a DNN and train nonlinear DNN-based controllers with strict stability guarantees. The resulting controller was purely learning-based, but it might be possible to adapt this novel training method for the purpose of approximating MPC with strict guarantees.